# aoas/sm_enkf

### v2

@@ 5.1 Lorenz-96 model
Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method and the three HD-EnKFs equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, under different forcing $F$ at moderate dimension $p=40$ and high dimension $p=100$ with different $\sigma_{0}$ and $n$. It shows that the HD-EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) models. Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states that were much closer to the oracle than the inflation method although both strains of methods showed increased RMSEs with respect to $|F-8|$. The standard EnKF with the sample covariance estimator to $\boldsymbol{\Sigma}_{t}^{f}$ had much larger assimilation errors in all cases.
Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), the advantages of the proposed HD-EnKFs over the inflation method were more apparent for the higher dimension $p=100$. The analysis RMSEs of the HD-EnKFs kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF. This means that the HD-EnKFs are quite attractive as the geophysical models in earth science are of ultra high dimension. The figure also displays the percentages of divergence of the five methods. It shows that the inflation method endured quite high divergences which increased substantially with the model error (e.g. reached 50% when $\sigma=0.5$), even higher than the standard EnKF. In contrast, the three HD-EnKF methods had quite low divergence rates for all cases.
@@ 5.2 Shallow Water Equation model
where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~{}m}\mathrm{~{}s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~{}s}^{-1}$, $k$ is the diffusion coefficient specified as $5\times 10^{4}\mathrm{~{}m}^{2}\mathrm{~{}s}^{-1}$, and $L$ and $D$ are the domain limits set as $500\mathrm{~{}km}$ and $300\mathrm{~{}km}$, respectively. The SWE was discretized with a uniform grid spacing of $10\mathrm{~{}km}$ in the $z_{1}$ and $z_{2}$ directions and solved by the Lax-Wendroff method with a time step of $30$ seconds. Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ of dimension $p=3p^{\prime}=4650$. The initial height (depth) field was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~{}m},5.5\mathrm{~{}m}$ and $3.325\mathrm{~{}m}$, respectively. The initial velocity field was derived from the initial height field with the geostrophic relation.
@@
Figure 3: RMSEs to the oracle analysis state (the left panels) and the true state (the right panels) at each assimilation time step by the standard, inflation, banding, tapering and threshold under the correct model ($k=5\times 10^{4}$) and imperfect model ($k=10^{4}$) with $n=100$, $p=4650$ and $q=930$ on the SWE model.
Table 2: Average RMSEs to the oracle analysis state and true state over the last 24 hours (2880 steps) of the simulation experiments and their $25\%$ and $75\%$ quantiles (in parentheses) by the five EnKF schemes for the SWE model under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models with $n=100$, and $p=4650$ and $q=930$.
@@
Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods (the standard EnKF, inflation and three HD-EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models while Table 2 summarizes the average analysis RMSEs over the last 24 hours (2880 steps), together with their $25\%$ and $75\%$ quantiles. Figure 3 shows the three proposed HD-EnKF methods outperformed the inflation and the standard EnKF methods for both the correct and imperfect models with much smaller RMSEs after the system settled down and that the superiority became more apparent over time, especially under the misspecified model. Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and slight increases RMSEs for the inflation method under true models.
@@
It is obvious from Table 2 that the HD-EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands with more evident superiority under imperfect models. Specifically, the proposed HD-EnKFs were able to reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated outcomes even under extremely sparse observations and ultra-high dimensional state variables. Thus, the proposed methods are highly appealing, especially considering that geophysical models in earth science often possess ultra-high dimensions.
6 Sea Temperature Assimilation in Northwest Pacific
This section demonstrates how the proposed method is applied to sea temperature assimilation based on the oceanography model named Regional Oceanic Modeling System (ROMS) with satellite data. ROMS is a three-dimensional, free-surface, terrain-following ocean general circulation model that solves the Reynolds-averaged Navier–Stokes equations using the hydrostatic and Boussinesq assumptions (Shchepetkin and McWilliams, 2005; Haidvogel et al., 2008). While the satellites produce high-resolution sea surface temperature data, they can not provide ocean measurement beneath the ocean surface where the physical model ROMS plays a key role.
The data assimilation was conducted from January 1st, 2008 to March 31st, 2008 over a study area 25.5N-34.7N, 144.4E-154E. The spatial resolution of the ROMS was $0.56^{\circ}\times 0.49^{\circ}$ with $18\times 20$ horizontal grids and $18$ uneven vertical layers over the water column, making the dimension $p$ for the state variable of sea temperature ($\mathbf{x}_{t}$) to be $6480=18\times 20\times 18$. To scale down the computation, we did not include the other ocean variables namely salinity and current velocity as the assimilated variables, and let them be internally determined by ROMS at each step with the sea temperature as the assimilated variables. The assimilation time length was daily which was also the frequency of the observations. The satellite data were the daily Optimum Interpolation Sea Surface Temperature (OISST; https://www.ncei.noaa.gov/products). OISST is a long-term data records for Sea Surface Temperature (SST) that incorporate satellite observations into regular global grids using ships and buoys observations correct satellite SST biases (Banzon and Reynolds, 2013). As the OISST only provides sea surface temperature, the observation dimension was $q=360$.
To run ROMS, we need to input the initial state of sea temperature $\mathbf{x}_{1}$, the boundary conditions and smoothed bathymetry data (sea bottom topography). The bathymetry data came from ETOPO2 (https://www.ncei.noaa.gov/products/etopo-global-relief-model), which is a global map of the earth’s seafloor and land elevations as shown in Figure S.1 of the SM. The initial state ($\mathbf{x}_{1}$) was from the Global Ocean Reanalysis and Simulations (GLORYS; https://data.marine.copernicus.eu/products) on January 1st. Thorough descriptions of boundary conditions of ROMS can be found in Marchesiello et al. (2001).
As the ocean temperature, salinity and current fields have different spatial scales and resolution in the three dimension study region both in the ROMS and the physical ocean, we transform the study region to a scale down space to facilitate the bandwidth selection in the banding estimator. Specifically, the grid points in the study region are denoted by coordinates $\{(u,v,h),u=1,\cdots 18,v=1,\cdots 20,h=1,\cdots 18\}$ in the transformed space of unit rectangles, where the banding estimator is performed. The temperatures on the physical grids are labelled as $\{x_{u,v,h}\}$ in the standardized grids and the whole state vector can be labelled by $\mathbf{x}=(x_{1,1,1},\cdots,x_{18,20,18})^{\mathrm{T}}$ with the subscript $t$ hidden. The banding estimator with bandwidth $k$ means that we keep the covariances between $x_{u,v,h}$ and $x_{u^{\prime},{v^{\prime}},h^{\prime}}$ on the standardized grids if they satisfy $\max\{|u^{\prime}-u|,|v^{\prime}-v|,|h^{\prime}-h|\}\leq k$.
We performed the HD-EnKF using the banding estimator with an ensemble size of $n=30$ to obtain the assimilated three-dimensional sea temperature for the study period and region. We also compared the proposed method with the standard EnKF and the inflation method. The out-of-sample performances of these methods were assessed on data collected under the ocean Argo project over the study period and region, accessed at https://argo.ucsd.edu/data/. The Argo project collects information from inside the ocean using floats that initially drift with the ocean currents and then surface up and down between the surface and mid-sea to collect measurements. As the observations provided by Argo are in-situ observations at various depths, they are more precise than the remote sensing data by the satellites, and hence, are used as substitutes for the ground truth for the out-of-sample assessment. The bandwidth $k$ used in the banding estimator was $k=1$, selected by minimizing the out-of-sample RMSE over the first month of the study period between assimilated temperatures and temperatures recorded by Argo at different depths. The inflation factor for the inflation method was similarly selected, which was $1.1$.
Figure 4: Daily assimilation RMSEs and correlations to the OISST(left panels) and the Argo data (right panels) from Jan. 1st 2008 to Mar. 31st 2008 using the standard EnKF (blue), inflation (red) and HD-EnKF with banding (green), respectively. The dashed horizontal line represents the average assimilation RMSEs and correlations over the whole assimilation period.
Table 3: The average assimilation RMSEs and correlations and their standard errors (in the parentheses) to the OISST data and the Argo data over the whole assimilation period.
Method OISST Argo
RMSE Corr RMSE Corr
standard 0.677(0.0067) 0.943(0.0012) 1.242(0.0421) 0.984(0.0013)
inflation 0.647(0.0071) 0.948(0.0011) 1.244(0.0421) 0.983(0.0012)
banding 0.277(0.0043) 0.990(0.0003) 1.134(0.0425) 0.986(0.0014)
Figure 5: The spatial distribution of the assimilated sea surface temperature on January 1st, February 15th and March 31st in 2008 over the study area, with the fill color from red to blue representing sea surface temperature from high to low.
Figure 4 displays both the in-sample RMSEs and correlations between the assimilated sea surface temperature and the OISST observations, and the out-of-sample RMSEs and correlation between the assimilated sea water temperature and the temperatures recorded by Argos at various depths. From Figure 4, the HD-EnKF provided much lower and more stable assimilation RMSEs and much higher correlations in the entire assimilation period than those for the standard EnKF and the inflation methods. On the out-of-sample performance on the Argo data, the HD-EnKF produced assimilated sea temperatures with smaller RMSEs and higher correlations both on average and for most days during the assimilation period. These imply that the proposed HD-EnKF algorithm could offer more accurate and robust assimilation results in practice. Detailed numerical results of these methods are reported in Table 3, which summarizes the average assimilation RMSEs and correlations of the assimilated $\hat{\mathbf{x}}_{t}^{a}$ to the OISST and the Argo data over the whole assimilation period. The spatial distribution of the assimilated sea surface temperature over the study area on January 1st, February 15th and March 31st in 2008 is depicted in Figure 5, which varied over time and saw an overall downward trend during the study period.
@@ 7 Discussion
In this paper, we propose HD-EnKF algorithms to suit the needs of high dimensional data assimilation. We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional cases for both the correct and misspecified models. The study shows that the proposed HD-EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations. One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies and the empirical study on sea temperature assimilation. The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive. The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are ultra-high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.

### v3

@@ 5.1 Lorenz-96 model
Figure 2 displays the average RMSEs to the oracle analysis state of the five assimilation methods, namely the standard EnKF, the inflation method and the three HD-EnKFs equipped with the banding and tapering (both respect the circular bandable structures), and the thresholding estimators, under different forcing $F$ at moderate dimension $p=40$ and high dimension $p=100$ with different $\sigma_{0}$ and $n$. It shows that the HD-EnKF with the three high dimensional covariance estimators had much smaller RMSEs than the standard EnKF and the inflation methods at both the correct ($F=8$) and misspecified ($F\neq 8$) models. Even when $F$ deviated more from the true value $8$, the HD-EnKFs still achieved analysis states much closer to the oracle than the inflation method, although both strains of methods showed increased RMSEs with respect to |F8||F-8|. The standard EnKF with the sample covariance estimator had much larger assimilation errors in all cases.
Comparing Figure 2’s panels (d)-(f) with panels (a)-(c), the advantages of the proposed HD-EnKFs over the inflation method were more apparent for the higher dimension $p=100$. The analysis RMSEs of the HD-EnKFs kept at a lower level especially when the force $F$ was equal or close to the true value ($F=8$), while the analysis RMSE of the inflation method ranked behind but better than the standard EnKF. This means that the HD-EnKFs are quite attractive as the geophysical models in earth science are of ultra high dimension. The figure also displays the percentages of divergence of the five methods. The inflation method endured quite high divergences which increased substantially with the model error (e.g. reached 50% when $\sigma=0.5$), even higher than the standard EnKF. In contrast, the three HD-EnKF methods had quite low divergence rates for all cases.
@@ 5.2 Shallow Water Equation model
where the fluid velocity components $u$ in $z_{1}$ direction and $v$ in $z_{2}$ direction as well as the fluid depth $h$ are the model variables, $g$ is the gravity acceleration $9.8\mathrm{~m}\mathrm{~s}^{-2}$, $f$ is the Coriolis parameter defined as constant $10^{-4}\mathrm{~s}^{-1}$, $k$ is the diffusion coefficient specified as $5\times 10^{4}\mathrm{~m}^{2}\mathrm{~s}^{-1}$, and $L$ and $D$ are the domain limits set as $500\mathrm{~km}$ and $300\mathrm{~km}$, respectively. The SWE was discretized with a uniform grid spacing of $10\mathrm{~km}$ in the $z_{1}$ and $z_{2}$ directions and solved by the Lax-Wendroff method with a time step of $30$ seconds. Thus, the SWE model had a dimension of $p^{\prime}=50\times 31$ for each state variable $u,v,h$, leading to the state vector $\mathbf{x}_{t}=(u_{t},v_{t},h_{t})^{\mathrm{T}}$ of dimension $p=3p^{\prime}=4650$. The initial height (depth) was given by $h(z_{1},z_{2})=H_{0}+H_{1}\tanh\big{(}{9(D/2-z_{2})}/2/D\big{)}+H_{2}\operatorname{sech}^{2}\big{(}9(D/2-z_{2})/{D}\big{)}\sin\big{(}{2\pi}z_{1}/{L}\big{)},$ where $H_{0},H_{1}$ and $H_{2}$ were set to be $50\mathrm{~m},5.5\mathrm{~m}$ and $3.325\mathrm{~m}$, respectively. The initial velocity field was derived from the initial height field with the geostrophic relation.
@@
Figure 3: RMSEs to the oracle analysis state (left panels) and the true state (right panels) at each assimilation step by the standard, inflation, banding, tapering and threshold under the correct model ($k=5\times 10^{4}$) and imperfect model ($k=10^{4}$) with $n=100$, $p=4650$ and $q=930$ on the SWE model.
Table 2: Average RMSEs to the oracle analysis state and true state over the last 24 hours (2880 steps) and their $25\%$ and $75\%$ quantiles (in parentheses) by the five EnKF schemes for the SWE model under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models with $n=100$, and $p=4650$ and $q=930$.
@@
Figure 3 reports the RMSEs at each assimilation time step of the five EnKF methods (the standard EnKF, inflation and three HD-EnKF) to the oracle analysis states $\mathbf{x}_{t}^{a}$ and the true states $\mathbf{x}_{t}$ under the correct ($k=5\times 10^{4}$) and the imperfect ($k=10^{4}$) models while Table 2 summarizes the average analysis RMSEs over the last 24 hours (2880 steps), together with their $25\%$ and $75\%$ quantiles. Figure 3 shows the three proposed HD-EnKF methods outperformed the inflation and the standard EnKF methods for both the correct and imperfect models with much smaller RMSEs after the system settled down and that the superiority became more apparent over time, especially under the misspecified model. Specifically, the RMSEs at each assimilation time step of the three proposed HD-EnKF schemes became lower and lower over time compared to increased RMSEs for the standard EnKF for all cases and for the inflation method under true models.
@@
It is obvious from Table 2 that the HD-EnKF methods had significantly lower RMSEs and narrower 25%-75% quantile bands with more evident superiority under imperfect models. Specifically, the proposed HD-EnKFs could reduce the RMSE by 41% compared to the inflation method, implying that the proposed methods could attain more accurate and robust assimilated results even under extremely sparse observations and high dimensional states. Thus, the proposed methods are highly appealing, especially considering the high dimensional geophysical models in earth science.
@@ 6 Discussion
In this paper, we propose HD-EnKF algorithms to suit the needs of high dimensional data assimilation. We take the opportunity to study the theoretical properties of the EnKF under both the fixed and the high dimensional cases for both the correct and misspecified models. The study shows that the proposed HD-EnKF can replace the widely used inflation method for data assimilation of a large dynamic system based on two considerations. One is that the established theoretical analysis on the consistency and the error bounds makes the proposed HD-EnKF methods more understandable with a theoretical guarantee; and the other is the better algorithm convergence and smaller RMSEs than the inflation method as shown in the simulation studies. The latter means that a smaller ensemble size for the EnKF can be used at a given level of accuracy which is particularly attractive as the running of the geophysical models is generally very expensive. The proposed HD-EnKF has great potential for applications in earth science where the geophysical models, the state variables or the observations are high dimensional, such as global carbon calculation, ocean data assimilation, weather forecasting, sea ice prediction and the generation of high-resolution data sets.